A gas held at 2.20 liters is 755 mmHg @ 59.8 degrees Celsius. What is the pressure if the temperature is raised to 298 degrees Celsius?
Answer #1

Using ideal gas law,

P1/T1 =P2/T2

755/332.8 =P2/571

P2 =1295.4 mm Hg

So pressure is 1295.4 mmHg